OpenStudy (anonymous):

Evaluate the integral using the properties of even and odd functions as an aid.

OpenStudy (anonymous):

\[\int\limits_{-2}^{2} x ^{2}+1)dx\]

OpenStudy (inkyvoyd):

\(\Huge =2\int_0^2 (x^2+1)\) because x^2+1 is an even function

OpenStudy (inkyvoyd):

I forgot the dx

OpenStudy (anonymous):

alright I know the answer for this is 272/15 but how do you do this Im new to this

OpenStudy (inkyvoyd):

well, it depends. there are a few ways to find this integral (power rule, limits, etc), and to know how to help you I need to know what course you are in. Are you currently in a pre-calc class covering basics of calculus, or is this a fully fledged calculus class?

OpenStudy (anonymous):

No i'm in Calculus AB

OpenStudy (inkyvoyd):

the integral of x^2+1 is x^3/3+x+c

OpenStudy (anonymous):

What confused is how come there is 3?

OpenStudy (anonymous):

do you add the power?

OpenStudy (anonymous):

by 1

OpenStudy (inkyvoyd):

Dude, shouldn't you be in the middle of the year of a calculus AB class?

OpenStudy (anonymous):

Yea!

OpenStudy (inkyvoyd):

I'm pretty sure you should've learned this a while back...

OpenStudy (anonymous):

My teacher is making me confused... he just recently!

OpenStudy (inkyvoyd):

the integral of \(x^n\) is \(\large \frac{x^{n+1}}{n}\)

OpenStudy (inkyvoyd):

+C of course

OpenStudy (anonymous):

Yea I know that part!

OpenStudy (inkyvoyd):

that is a basic rule you should know. Apply it to what you have.

OpenStudy (anonymous):

I'm confused with even and odd that's all

OpenStudy (inkyvoyd):

what's the integral of x^3+1?

OpenStudy (inkyvoyd):

do you know what an even and odd function is?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Can you explain to me please?

OpenStudy (inkyvoyd):

an odd function is a function such that f(-x)=-f(x). An even function is a function such that f(-x)=f(x). Graphically an even function is symmetrical about the y-axis while an odd function is symmetrical when rotated about the origin 180 degrees. Sin(x) is an odd function while cos(x) is an even function. using the laws of definite integrals, you can break up the integrals. If you know a function is odd, and say have to find the deifnite integral from -10 to 10, you know it is the same thing as finding the integral from 0 to 10 and multiplying by two.

OpenStudy (anonymous):

THANKS!!!

OpenStudy (inkyvoyd):

read http://www.purplemath.com/modules/fcnnot3.htm for more information. Note that because you can "fold" the first picture they gave you in half (it's an even function), you only have to find the area of the right half of the parabola to find the area of the left half of the parabola

OpenStudy (anonymous):

Oh Thanks! :)

OpenStudy (anonymous):

Now I remember what my teacher means to say.